### v1

# Sublinear hitting sets for some geometric graphs

Abstract
For an $n$-vertex graph $G$, let $h(G)$ denote the smallest size of a subset of $V(G)$ such that it intersects every maximum independent set of $G$. A conjecture posed by Bollobás, Erdős and Tuza in early 90s remains widely open, asserting that for any $n$-vertex graph $G$, if the independence number $\alpha(G)=\Omega(n)$, then $h(G)=o(n)$. In this paper, we establish the validity of this conjecture for various classes of graphs, including disk graphs, even-hole-free graphs, circle graphs, and those hereditary graphs having sublinear balanced separators. We also determine the exact values of smallest possible hitting sets in comparability graphs, incomparability graphs and the graphs with VC-dimension one.
1 Introduction
1.1 Overview
For a graph $G=(V,E)$, an independent set is a set of vertices in $G$, no two of which are adjacent. A maximum independent set is an independent set of the largest possible size for the graph $G$, and this size is called the independence number of $G$ and is usually denoted by $\alpha(G)$. We say $T\subseteq V(G)$ is a hitting set of the graph $G$ if $T\cap I\neq\emptyset$ holds for every maximum independent set $I\subseteq V(G)$. Let $h(G)$ be the smallest size of a hitting set $T\subseteq V(G)$. Assume that $n$ is sufficiently large whenever this is needed. Bollobás, Erdős and Tuza (see, [13, 20]) raised the following conjecture.
Conjecture 1 ([13, 20]).
Let $G$ be an $n$-vertex graph with $\alpha(G)=\Omega(n)$, then $h(G)=o(n)$.
For this conjecture, there was a clever observation of Hajnal [36], (also see Lemma 1 in [49]) which states that for all maximum independent sets in $G$, say $A_{1},A_{2},\ldots,A_{s}$, the following inequality always holds:
$|A_{1}\cap A_{2}\cap\cdots\cap A_{s}|+|A_{1}\cup A_{2}\cup\cdots\cup A_{s}|\geq 2\alpha(G).$
As a simple consequence, one can see that if $\alpha(G)>\frac{|G|}{2}$, then $|A_{1}\cap A_{2}\cap\cdots\cap A_{s}|>0$, which implies that $h(G)=1$. However, 1 even remains open in the case of $c=\frac{1}{2}-\varepsilon$ for any small $\varepsilon>0$. A recent major progress of 1 due to Alon [3] showed that for every $n$-vertex regular graph $G$ with $\alpha(G)>(\frac{1}{4}+\varepsilon)n$ for arbitrarily small $\varepsilon>0$, we have $h(G)\leq O(\sqrt{n\log{n}})$. Alon’s proof elegantly combined the container method [5, 6, 50] and the above observation of Hajnal. On the other hand, Alon [3] provided an explicit construction of graph $G$ with $\alpha(G)=\Theta(n)$ and $h(G)=\Theta(\sqrt{n})$, which means that the $o(n)$ term in 1 cannot be replaced by $o(\sqrt{n})$. Moreover, Alon provided another type of construction of graph $G$ with $n=2^{m}$ vertices such that $\alpha(G)=\sum_{i=0}^{\frac{m}{2}-t}\binom{n}{i}$ and $h(G)=\Theta(m^{2})$. This construction is mainly based on some ideas from discrepancy theory and Kleitman’s isodiametric inequality [41], which also disproves a conjecture of Dong and Wu [17]. In particular, Alon [2] suggested to study whether 3-colorable graphs satisfy 1, and showed that for every integer $q\geq 1$, there exists an $8$-colorable graph $G_{q}$ with $h(G)>q$, see [35, Theorem 1.3].
Notice that the condition “maximum independent sets” cannot be replaced by “maximal independent sets”. To see this, let $G$ be a book graph that consists of a clique $K_{(1-c)n}$, and the other $cn$ vertices are only adjacent to all vertices in the clique $K_{(1-c)n}$. Then we need exactly $(1-c)n+1$ vertices to hit all maximal independent sets.
One way to better understand 1 is to test whether it holds for specific classes of graphs. In recent work of Hajebi, Li and Spirkl [35], the authors put this idea into practice. More precisely, they proved that for any induced $P_{5}$-free graph $G$, $h(G)\leq f(\omega(G))$, where $P_{5}$ is a path with $5$ vertices, $f:\mathbb{N}\rightarrow\mathbb{N}$ is a function and $\omega(G)$ is the largest size of a clique in $G$. A family of graphs $\mathcal{G}$ is termed hereditary if, for any graph $G\in\mathcal{G}$ and any induced subgraph $G^{\prime}$ of $G$, we have $G^{\prime}\in\mathcal{G}$. We say that a hereditary class $\mathcal{G}$ of graphs has the Erdős-Hajnal property if there exists some $\varepsilon>0$ such that every graph $G\in\mathcal{G}$ satisfies either $\alpha(G)>|V(G)|^{\varepsilon}$ or $\omega(G)>|V(G)|^{\varepsilon}$. The famous Erdős-Hajnal conjecture [21] states that every hereditary class of graphs has Erdős-Hajnal property. Hajebi, Li and Spirkl [35] recently provided a new approach to this conjecture. More precisely, they showed that for a hereditary class $\mathcal{G}$ of graphs, if every graph $G\in\mathcal{G}$ satisfies $h(G)\leq\textup{Poly}(\omega(G))$, then $\mathcal{G}$ satisfies Erdős-Hajnal property.
However, this approach does not always work. Note that the family of graphs with VC-dimension at most $d\in\mathbb{N}$ is hereditary.
Theorem 1.1.
There exists some graph $G$ with VC-dimension at most $3$ and $\omega(G)=3$ such that $h(G)\geq\frac{\sqrt{|G|}}{2}$.
Theorem 1.1 follows from Alon’s construction in [3, Theorem 1.3], we will discuss this in Section 5.4.1.
1.2 Our contributions
In this paper, we will demonstrate the validity of 1 for several important classes of graphs, some of which stem from geometry, while others exhibit highly favorable combinatorial properties.
A hole in a graph is an induced subgraph which is a cycle of length at least four. A hole is called even if it has an even number of vertices. An even-hole-free graph is a graph with no even holes. One of the most important sub-classes of the even-hole-free graphs is the family of chordal graphs. A chordal graph is a simple graph in which every cycle of length four and greater has a cycle chord.
The absence of even-hole-free graphs attracts broad attention in the field of graph theory. The renowned strong perfect graph theorem, established by Chudnovsky, Robertson, Seymour, and Thomas [11], states that a graph achieves perfection precisely when it forbids both odd holes and odd anti-holes (defined as the complements of holes). Graphs without any odd holes and odd anti-holes are termed Berge graphs. It has been observed that the structure of even-hole-free graphs bears striking similarities to that of Berge graphs, for instance, the prohibition of induced $4$-holes necessitates the exclusion of any anti-hole with a length of at least $6$. For a positive integer $\gamma$, we define a vertex $v\in V(G)$ to be $\gamma$-simplicial if its neighborhood $N_{G}(v)$ forms a disjoint union of at most $\gamma$ cliques. Chudnovsky and Seymour [12] recently provided a valid proof and showed that every even-hole-free graph contains a $2$-simplicial vertex, thereby settling a conjecture posed by Reed [38]. It then follows that $h(G)\leq 2\omega(G)+1$ by 2.1. We also refer interested readers to a survey for even-hole-free graphs [56]. Here we confirm that 1 holds for any even-hole-free graph.
Theorem 1.2.
Let $G$ be an even-hole-free graph with $\alpha(G)=cn$ for some $c>0$, then we have $h(G)\leq\frac{5n}{\log{n}}$.
To prove Theorem 1.2, we develop a framework to show that graphs with certain locally sparse properties satisfy 1, see Theorem 3.1. Under this framework, we can also prove 1 holds for the disk graph, which is the intersection graph of disks in two-dimensional space.
Theorem 1.3.
Let $G$ be an $n$-vertex disk graph with $\alpha(G)=cn$ for some $c>0$, then we have $h(G)\leq\frac{25n}{\log{n}}$.
Our subsequent result concerning disk graphs follows in a similar vein as the recent work of Chudnovsky and Seymour [12], which is also related to the problem of Hajebi, Li and Spirkl [35].
Theorem 1.4.
Every disk graph $G$ contains an $11$-simplicial vertex. Consequently, every disk graph $G$ satisfies $h(G)\leq 11\omega(G)+1$.
The constant $11$ above might be improved and we make no attempt to optimize it here. There are some special sub-families of the disk graphs, for instance, a unit disk graph is a disk graph with an intersection graph consisting of disks of unit size, one can easily show $h(G)=O(1)$ when $\alpha(G)=cn$ since unit disk graph $G$ is induced $K_{1,6}$-free. A coin graph is a disk graph with the additional condition that no two disks have more than one point in common. Note that the circle packing theorem, first proven by Paul Koebe [42] in 1936, states that a graph is planar only if it is a coin graph. Therefore any $n$-vertex coin graph has at most $63n-$ 6 edges, which implies we can only use $6$ vertices to hit all maximum independent sets in any coin graph by 2.1.
Moreover, the results in Theorem 1.3 can be generalized to the intersection graph of higher dimensional balls by simple modification of our proof. One of the useful auxiliary results is that, in any $d$-dimensional space, the number of unit balls that can touch a unit ball is bounded by $f(d)$ for some function $f$ [48]. For instance, $f(2)=6$ implies that the unit disk graph is induced $K_{1,6}$-free.
Next, our attention shifts to the circle graph $G$, where $V(G)$ comprises a finite set of chords on a circle, with two vertices being adjacent if and only if their corresponding chords intersect. Circle graphs and their representations play a pivotal role in numerous unexpected fields. Exemplary instances include knot theory [7] and quantum computing [54]. Also, circle graphs have garnered considerable interest both in graph theory [8, 31, 32] and computer science [9, 39, 40]. In particular, a very recent breakthrough of Davies and McCarty [15] showed that the chromatic number of circle graphs is quadratic in its clique number, which significantly improves the previous results in [43, 44]. Shortly after this, the upper bound was further improved by Davies [14], who showed the sharp upper bound $\chi(G)\leq O(\omega(G)\log{\omega(G)})$ when $G$ is a circle graph. Further exploration of circle graphs can be found in an informative textbook [34]. We can show that 1 also holds for circle graphs.
Theorem 1.5.
Let $G$ be a circle graph with $\alpha(G)=cn$ for some $c>0$, then $h(G)\leq\frac{2(1-c)n\log\log{n}}{c\log{n}}+1$.
Given a partially ordered set $(P,\prec)$, its comparability graph is a graph with vertex set corresponding to elements of $P$ with two vertices sharing an edge if and only if the corresponding elements are comparable in $P$. On the other hand, an incomparability graph is a graph that connects pairs of elements of a poset that are not comparable to each other in a partial order, i.e. it is the complement of a comparability graph. The graph on posets is one of the important topics in the field of extremal combinatorics, see [22, 29]. For both the comparability graphs and the incomparability graphs, we can prove the sharp results on $h(G)$.
Proposition 1.6.
Let $G$ be an $n$-vertex comparability graph, then $h(G)\leq\frac{n}{\alpha(G)}$. In particular, if $\alpha(G)=cn$ for some $c>0$, then $h(G)\leq\lfloor\frac{1}{c}\rfloor$.
Proposition 1.7.
Let $G$ be an $n$-vertex incomparability graph, then $h(G)\leq\frac{n}{\alpha(G)}$. In particular, if $\alpha(G)=cn$ for some $c>0$, then $h(G)\leq\lfloor\frac{1}{c}\rfloor$.
Assume that $\alpha(G)\mid n$, there are two types of constructions attaining the upper bounds on the above theorems. The first one the complete $\frac{n}{\alpha(G)}$-partite graph, the second one is the disjoint union of $\alpha(G)$ many cliques of size $\frac{n}{\alpha(G)}$. We will provide a short proof of 1.6 via simply applying Dilworth’s theorem, and we also give an alternative proof in [10], which might be useful in other graphs.
Let $\mathcal{F}\subseteq 2^{[n]}$ be a set system, the Vapnik-Chervonenkis dimension (for short, VC-dimension) of $\mathcal{F}$ is the largest integer $d$ for which there exists a subset $S\subseteq[n]$ with $|S|=d$ such that for every subset $B\subseteq S$, one can find a member $A\in\mathcal{F}$ with $A\cap S=B$. We say that the graph $G$ has VC-dimension $d$ if the set system :=$\mathcal{F}:=\{N_{G}(v):v\in V(G)\}$ induced by the neighborhood of vertices in $G$ has VC-dimension $d$. The VC-dimension is one of the most useful combinatorial parameters that measure its complexity. The notion was introduced by Vapnik and Chervonenkis [55] in 1971, as a tool in mathematical statistics.
We can derive the sharp upper bound on $h(G)$ for graphs with VC-dimension one.
Theorem 1.8.
Let $G$ be an $n$-vertex graph with VC-dimension $1$, then $h(G)\leq\max\{\frac{n}{\alpha(G)},3\}$. In particular, if $\alpha(G)=cn$ for positive constant $c\neq\frac{2}{5}$, then $h(G)\leq\lfloor\frac{1}{c}\rfloor$; if $\alpha(G)=\frac{2n}{5}$, then $h(G)\leq 3$.
The upper bounds on Theorem 1.8 are sharp for different region of $c$. The exact results reflect a rare phenomenon, where the performance of $h(G)$ is highly sensitive to the selection of $c$. Indeed, when $c\neq\frac{2}{5}$, if $\lfloor\frac{1}{c}\rfloor$ is an integer, one can take the complete balanced $\frac{1}{c}$-partite graph, otherwise, one can take complete $\lceil\frac{1}{c}\rceil$-partite graph with parts $A_{1},A_{2},\ldots,A_{\lceil\frac{1}{c}\rceil}$ such that $|A_{1}|=(1-c\cdot\lfloor\frac{1}{c}\rfloor)n$ and $|A_{i}|=cn$ for any $i\geq 2$. It is easy to check that the complete multipartite graph has VC-dimension one. Particularly when $c=\frac{2}{5}$, we can see $G=C_{5}[\frac{n}{5}]$ satisfies $h(G)=3$ and the VC-dimension of $G$ is one. Note that when $\frac{1}{3}<c\leq\frac{1}{2}$, only when $\frac{2}{5}$, $h(G)$ might be $3$, while in other cases $h(G)$ must be strictly less than $3$.
For convenience, we construct a partially ordered diagram illustrated in Fig. 1.1 to depict the principal findings of this paper and elucidate the interconnections among various graph classes, as well as some open problems.
Circle $k$-gon graph: $O(\frac{n\log{\log{n}}}{\log{n}})$Interval filament graph: 4Even-hole-free graph: $O(\frac{n}{\log{n}})$String graph: 3Chordal graph: $O(\frac{n}{\log{n}})$Circle graph: $O(\frac{n\log{\log{n}}}{\log{n}})$Trapezoid graph: $\frac{n}{\alpha(G)}$ncomparability graph: $\frac{n}{\alpha(G)}$ircular-arc graph: $O(\frac{n}{\alpha(G)})$Interval graph: $\frac{n}{\alpha(G)}$omparability graph: $\frac{n}{\alpha(G)}$isk graph: $O(\frac{n}{\log{n}})$Graph with bounded VC-dimension: 2Unit disk graph: $O(\frac{n}{\alpha(G)})$Graph with VC-dimension one: $\frac{n}{\alpha(G)}$
Figure 1.1: Upper bounds on $h(G)$ for various families of graphs $G$ with $\alpha(G)=\Omega(n)$. $A\rightarrow B$: $B$ is a sub-family of $A$
Structure of this paper. In Section 2, we introduce some basics on this topic and We provide the proofs of 1.61.7 and Theorem 1.8. In Section 3, we will show more general results for graphs with certain locally sparse properties in Theorem 3.1. We also discuss the combinatorial properties of disk graphs and even-hole-free graphs and provide the proofs of Theorem 1.2 and Theorem 1.3 in detail. We will prove Theorem 1.5 in Section 4. We conclude this paper and propose several conjectures related to the main topic of this paper in Section 5.
Notations. For a vertex $u\in V(G)$, we will use $N_{G}(u)$ (or $N(u)$ if the subscript is clear) to denote the set of the neighborhood of $u$. Let $\delta(G):=\min_{v\in V(G)}|N_{G}(v)|$ be the minimum degree of $G$. Let $N_{G}(A)=\bigcup_{a\in A}N_{G}(a)$. For two sets $A$ and $B$, let $A\setminus B:=\{x\in A,x\notin B\}$ and let $A\triangle B:=(A\setminus B)\cup(B\setminus A)$ be the symmetric difference between $A$ and $B$. For a graph $H$, we write $H[t]$ the $t$-blowup of $H$ by replacing every vertex of $H$ by an independent set of size $t$ and every edge of $H$ by a copy of $K_{t,t}$ and we write $H[\cdot]$ when mentioning a blow-up of $H$ without specifying its size (and the sizes of these independent sets could be different). We use $\mathcal{I}_{\textup{max}}(G)$ to denote the collection of all maximum independent sets in $G$. We say a graph $G$ is acyclic if $G$ does not contain any cycle.
The notations $o(\cdot)$, $O(\cdot)$, $\Theta(\cdot)$ and $\Omega(\cdot)$ have their usual asymptotic meanings. We use $\log$ throughout to denote the base $2$ logarithm. For the sake of clarity of presentation, we omit floors and ceilings and treat large numbers as integers whenever this does not affect the argument.
2 Warm-ups
2.1 Basics
Regarding the main problem in this paper, there are some simple results as follows.
Proposition 2.1.
  1. (1)
    If $G$ has connected components $G_{1},G_{2},\ldots,G_{m}$, then $h(G)\leq\min\limits_{i}h(G_{i})$.
  2. (2)
    For any vertex $v\in V(G)$, $\{v\}\cup N_{G}(v)$ is a hitting set of $G$. In particular, $h(G)\leq\delta(G)+1$.
  3. (3)
    If $G$ is a bipartite graph, then $h(G)\leq 2$.
Proof of 2.1.
For (1), note that every $I\in\mathcal{I}_{\textup{max}}(G)$ is a disjoint union of $I_{i}\in\mathcal{I}_{\textup{max}}(G_{i})$ for $i\in[m]$, therefore, to hit all maximum independent sets in $G$, it suffices to hit all maximum independent sets in a connected component. For (2), if there is some $I\in\mathcal{I}_{\textup{max}}(G)$ such that $I\cap N_{G}(v)=\emptyset$, then $v\in I$. For (3), let $G=A\cup B$, if $|A|\neq|B|$, then $\alpha(G)>\frac{|G|}{2}$, which yields $h(G)=1$ by Hajnal’s observation. Assume $|A|=|B|$, if there is a subset $S\subseteq A$ such that $|S|>|N(S)|$, then $\alpha(G)\geq|S\cup B\setminus N(S)|>\frac{|G|}{2}$, which implies $h(G)=1$. Otherwise, by Hall’s theorem [37], there is a perfect matching between $A$ and $B$. Take one edge $ab$ from this matching and put two endpoints $a$ and $b$ into the hitting set. It is easy to check that there is no independent set of size $\frac{|G|}{2}$ in the induced subgraph on subset $V(G)\setminus\{a,b\}$. ∎
2.2 Comparability graphs and incomparability graphs
An antichain in a poset $P$ is a set of elements such that no two of which are comparable to each other, and a chain is a set of elements such that every two of which are comparable. A chain decomposition is a partition of the elements of the order into disjoint chains. Dilworth’s theorem [16] states that, in any finite poset, the largest antichain has the same size as the smallest chain decomposition.
Theorem 2.2 (Dilworth [16]).
Let $P$ be a finite poset, then
$\min\{m:\leavevmode\nobreak\ \exists\leavevmode\nobreak\ \textup{chains\ }C_{1},C_{2},\ldots,C_{m}\leavevmode\nobreak\ \textup{s.t.}\leavevmode\nobreak\ P=\bigcup_{i=1}^{m}C_{i}\}=\max\{|A|:A\leavevmode\nobreak\ \textup{is\ an\ antichain}\}$
Proof of 1.6.
Let $P$ be a poset and $G$ be a comparability graph on $P$. Note that $\alpha(G)$ is the maximum size of antichain in $P$. By Theorem 2.2, there is a chain decomposition of $P$ of size $\alpha(G)$, namely $P=C_{1}\cup C_{2}\cup\cdots\cup C_{\alpha(G)}$. We claim that for each $i\in[\alpha(G)]$, $C_{i}$ is a hitting set for $\mathcal{I}_{\textup{max}}(G)$. Indeed, by Theorem 2.2 again, we see that the size of largest antichain in $P\setminus C_{i}$ is at most $\alpha(G)-1$.
Therefore by pigeonhole principle, there is a chain $C_{i}$ with size at most $\frac{n}{\alpha(G)}$ such that $C_{i}$ hits all maximum independent sets in $G$. ∎
Proof of 1.7.
Let $P$ be a poset and $G$ be the incomparability graph on $P$ with $\alpha(G)=\alpha$. For every maximum independent set $I$ in $G$, label its vertices as $b_{1}\prec\cdots\prec b_{\alpha}$, and color each vertex $b_{i}$ with color $i$.
Claim 2.3.
Every vertex in $G$ receives at most one color.
Proof of claim.
Suppose that there is a vertex $v$ receiving two colors. Thus, there are two chains of length $\alpha$ containing $v$, say $\mathcal{X}=x_{1}\prec\cdots\prec x_{\alpha}$ and $\mathcal{Y}=y_{1}\prec\cdots\prec y_{\alpha}$, such that $v=x_{i}=y_{j}$ with $i<j$. Then $y_{1}\prec y_{2}\prec\cdots\prec y_{j-1}\prec v\prec x_{i+1}\prec\cdots\prec x_{\alpha}$ is a chain of length $\alpha+1$, a contradiction. ∎
Therefore, we colored all vertices that belong to a maximum independent set with $\alpha$ colors. For any color $i\in[\alpha]$, the set of all vertices with color $i$ hits all maximum independent sets of $G$. Hence, by the pigeonhole principle, we have $h(G)\leq\frac{n}{\alpha}$. ∎
2.3 Graphs with VC-dimension $1$: Proof of Theorem 1.8
Let $G$ be a graph with VC-dimension $1$ and $\alpha(G)=\alpha$.
Lemma 2.4.
Let $G$ be a graph with VC-dimension one, then $G$ is induced $P_{5}$-free.
Proof of Lemma 2.4.
Suppose there is an induced copy of $P_{5}:=x_{1}x_{2}x_{3}x_{4}x_{5}$, then $\{x_{2},x_{4}\}$ is shattered, a contradiction. ∎
Moreover, we can also obtain the following simple structural property.
Lemma 2.5.
Let $G$ be a graph with VC-dimension one. If there is a triangle in $G$ with vertices $a_{1},a_{2},a_{3}$, then every vertex in $V(G)\setminus\{a_{1},a_{2},a_{3}\}$ is adjacent to at least two vertices in $\{a_{1},a_{2},a_{3}\}$.
Proof of Lemma 2.5.
If a vertex $v\in V(G)\setminus\{a_{1},a_{2},a_{3}\}$ is not adjacent to any vertex in $\{a_{1},a_{2},a_{3}\}$, then the set $\{a_{1},a_{2}\}$ is shattered. Moreover, if $v$ is adjacent to exactly one of $\{a_{1},a_{2},a_{3}\}$, by symmetry, assume $va_{1}\in E(G)$, then $\{a_{2},a_{3}\}$ is shattered. ∎
We will also take advantage of the structural property of triangle-free graphs without induced $P_{5}$, which was proven by Sumner [52] in 1981.
Theorem 2.6 ([52]).
Let $G$ be an induced $P_{5}$-free and triangle-free graph, then each connected component of $G$ is either a bipartite graph, or a copy of $C_{5}[\cdot]$.
Suppose that $G$ is triangle-free, then by Theorem 2.6, if $G$ is bipartite, then 2.1 yields that $h(G)\leq 2$, otherwise $G=C_{5}[\cdot]$ with parts $F_{1},F_{2},F_{3},F_{4},F_{5}$, and one can see $\alpha(G)\geq\frac{2n}{5}$. We further notice that if $c\neq\frac{2}{5}$, then $h(G)\leq 2$. To see this, if $c\neq\frac{2}{5}$, then there exists some pair $(i,j)$ with $|i-j|\not\equiv 1\pmod{5}$ such that $|F_{i}\cup F_{j}|<\frac{2n}{5}$. Without loss of generality, assume $|F_{1}\cup F_{3}|<\frac{2n}{5}$, then we can pick arbitrary one vertex $u\in F_{4}$ and $v\in F_{5}$ respectively. Obviously for every $I\in\mathcal{I}_{\textup{max}}(G)$, $I\cap\{u,v\}\neq\emptyset$, which yields $h(G)\leq 2$. when $c=\frac{2}{5}$, if $|F_{i}|=\frac{n}{5}$ for each $i\in[5]$, then $h(G)=3$, otherwise, we can see $h(G)\leq 2$ via the identical argument as above.
Next we can assume that there is a triangle with vertex set $\{a_{1},a_{2},a_{3}\}$. Take an arbitrary $I:=\{v_{1},v_{2},\ldots,v_{\alpha}\}\in\mathcal{I}_{\textup{max}}(G)$. Suppose that $I\cap\{a_{1},a_{2},a_{3}\}=\emptyset$, then by Lemma 2.5, $v_{1}$ must be adjacent to at least two vertices in $\{a_{1},a_{2},a_{3}\}$. Without loss of generality, we can assume that $v_{1}a_{1}\in E(G)$ and $v_{1}a_{2}\in E(G)$. As $v_{1},a_{1},a_{2}$ together form a copy of triangle, then for every vertex $v_{j}\in I\setminus\{v_{1}\}$, $v_{j}$ must be adjacent to both of $a_{1}$ and $a_{2}$ by Lemma 2.5, which indicates that all of vertices in $I$ are adjacent to both of $a_{1}$ and $a_{2}$. Moreover, as $a_{3}\notin I$, then $a_{3}$ must be adjacent to some vertex in $I$, say $a_{3}v_{k}\in E(G)$, which implies that $a_{1},a_{3},v_{k}$ form a copy of triangle. Again by Lemma 2.5, any vertex $v_{\ell}\in I\setminus\{v_{k}\}$ should be adjacent to $a_{3}$. Now let $v\in V(G)\setminus(I\cup\{a_{1},a_{2},a_{3}\})$. First, $v$ must be adjacent to at least two vertices in $\{a_{1},a_{2},a_{3}\}$, and we assume $a_{1}v,a_{2}v\in E(G)$ without loss of generality. As $v\notin I$, there must be some vertex $v_{q}$ such that $v_{q}v\in E(G)$, which implies that $a_{1}v_{q}v$ forms a copy of triangle, then by Lemma 2.5, every vertex in $I$ should be adjacent to the vertex $v$.
One can also apply the same argument to show that if $|I\cap\{a_{1},a_{2},a_{3}\}|=1$, every vertex $v\in V(G)\setminus I$ is adjacent to all vertices in $I$. We omit the repeated details here.
Therefore, if $G$ contains a copy of triangle as a subgraph, then all of the maximum independent sets of $G$ must be disjoint, which implies that $h(G)\leq\frac{n}{\alpha}$ in this case. This finishes the proof.
3 A general framework for locally sparse graphs
In this section, we provide a general framework to find a relatively small hitting set. This framework performs well for those graphs with linear independence number and certain locally sparse property. As examples, here we show that the even-hole-free graphs and disk graphs satisfy the property.
3.1 A general framework
Theorem 3.1.
Let $c>0$, $s\in\mathbb{N}$ and $G$ be an $n$-vertex graph with $\alpha(G)=cn$. If for any $I_{1},I_{2}\in\mathcal{I}_{\textup{max}}(G)$, $e(G[I_{1}\triangle I_{2}])\leq s\cdot|I_{1}\triangle I_{2}|$, then $h(G)\leq\frac{5sn}{\log{n}}$.
Proof of Theorem 3.1.
Let $\delta=\frac{a}{\log{n}}$ where $a=5s$ is a constant and let $b=0.98a$. Select a maximal sequence $I_{1},\ldots,I_{t}\in\mathcal{I}_{\textup{max}}(G)$ such that for each $k\in[t]$, $|I_{k}\setminus\big{(}\bigcup_{i=1}^{k-1}I_{i}\big{)}|>\frac{\delta n}{bk}$. Let $W=\bigcup_{i=1}^{t}I_{i}$ and $G^{\prime}=G[W]$. Then
$n\geq|W|\geq cn+\sum\limits_{i=2}^{t}\frac{\delta n}{bi}=cn+(\log{t}-O(1))\frac{\delta n}{b},$
and so $t\leq e^{b/\delta+O(1)}<\frac{n^{0.99}}{2s}$. Then, by the assumption, we have
$e_{G}(I_{1},\bigcup_{i=2}^{t}I_{i})\leq\sum_{i=2}^{t}e(G[I_{1}\triangle I_{i}])\leq 2s(t-1)cn<cn^{1.99}.$ (1)
Now suppose that $G^{\prime}=G$. By Eq. 1 and the pigeonhole principle, there exists a vertex $v\in I_{1}$ with $d_{G}(v)<n^{0.99}$ and we are done by 2.1.
It remains to consider the case that $G^{\prime}\neq G$. Again by Eq. 1, we can pick a subset $C\subseteq I_{1}$ with $|C|=\frac{\delta n}{4s(t-1)}$ such that $e_{G}(C,\bigcup_{i=2}^{t}I_{i})\leq 2s(t-1)|C|\leq\frac{\delta n}{2}$. Note that
$|C|+|N_{G^{\prime}}(C)|\leq|C|+e_{G}(C,\bigcup_{i=2}^{t}I_{i})\leq\delta n=\frac{5sn}{\log n}.$
It suffices to show that $C\cup N_{G^{\prime}}(C)$ hits all maximum independent sets in $G$. Fix an arbitrary $I\in\mathcal{I}_{\textup{max}}(G)$. The maximality of $t$ infers that $|I\setminus W|\leq\frac{\delta n}{b(t+1)}$. Suppose to the contrary that $I\cap(C\cup N_{G^{\prime}}(C))=\emptyset$. Then, $(I\cap W)\cup C$ is an independent set of size $|C|+|I|-|I\setminus W|\geq\frac{\delta n}{4s(t-1)}+(cn-\frac{\delta n}{b(t+1)})>cn$ as $b=0.99a>4s$, contradicting to $\alpha(G)=cn$. ∎
3.2 Even-hole-free graphs: Proof of Theorem 1.2
A well-known result states that the number of edges in an acyclic graph is less than the number of vertices. Based on Theorem 3.1, to prove Theorem 1.2, it suffices to show the following lemma.
Lemma 3.2.
Let $G$ be an even-hole-free graph. Then, for any $I_{1},I_{2}\in\mathcal{I}_{\textup{max}}(G)$, $G[I_{1}\triangle I_{2}]$ is acyclic.
Proof of Lemma 3.2.
Consider a minimal cycle $C$ in $G^{\prime}=G[I_{1}\triangle I_{2}]$ if exists. By minimality, $C$ is an induced cycle in $G$. As $G^{\prime}$ is bipartite, $C$ is an induced even cycle, a contradiction. Thus no such cycle exists and $G^{\prime}$ is acyclic as desired. ∎
3.3 Disk graphs: Proof of Theorem 1.3
Similarly as above, based on Theorem 3.1, to prove Theorem 1.3, it suffices to show the following lemma.
Lemma 3.3.
Let $G$ be an $n$-vertex disk graph with $\alpha(G)=cn$ for some $c>0$. For any $I_{1},I_{2}\in\mathcal{I}_{\textup{max}}(G)$, $e(G[I_{1}\triangle I_{2}])\leq 5|I_{1}\triangle I_{2}|$.
Proof of Lemma 3.3.
Firstly, we prove the following claim using the geometric intuition that a disk cannot intersect simultaneously with too many non-intersecting larger disks. For a vertex $v$, we denote $\rho(v)$ the radius of the disk corresponding to $v$.
Claim 3.4.
Let $v\in V(G)$ be a vertex with radius $\rho(v)$, then there is no independent set $I$ of size $6$ in the induced subgraph $G[N_{G}(v)]$ such that every vertex $u$ in $I$ satisfies $\rho(u)\geq\rho(v)$.
Proof of claim.
Suppose there are seven disks $A,B_{1},B_{2},\ldots,B_{6}$ with centers $\bm{x},\bm{y}_{1},\bm{y}_{2},\ldots,\bm{y}_{6}$ and radii $r_{A},r_{1},r_{2},\ldots,r_{6}$ such that for all $i\in[6]$, $r_{A}\leq r_{i}$, $A$ intersects $B_{i}$ and $B_{i}$ are pairwise disjoint. Then, for any $j\neq i$, $|\bm{x}-\bm{y}_{i}|\leq r_{A}+r_{i}\leq r_{j}+r_{i}<|\bm{y}_{i}-\bm{y}_{j}|$ and $|\bm{x}-\bm{y}_{j}|\leq r_{A}+r_{j}\leq r_{i}+r_{j}<|\bm{y}_{i}-\bm{y}_{j}|$. Therefore, the angle $\angle\bm{y}_{i}\bm{x}\bm{y}_{j}$ is greater than $\frac{\pi}{3}$ for any pair of $1\leq i<j\leq 6$. This yields $2\pi<\angle\bm{y}_{1}\bm{x}\bm{y}_{2}+\cdots+\angle\bm{y}_{5}\bm{x}\bm{y}_{6}+\angle\bm{y}_{6}\bm{x}\bm{y}_{1}=2\pi$, a contradiction. ∎
For a subset $S\subseteq V(G)$, let $d_{S}^{\geq}(u)$ be the number of neighbors $w\in N_{S}(u)$ such that $\rho(w)\geq\rho(u)$. For any $I_{1},I_{2}\in\mathcal{I}_{\textup{max}}(G)$, By 3.4 we have
$e(G[I_{1}\triangle I_{2}])\leq\sum\limits_{u\in I_{1}}d^{\geq}_{I_{2}\setminus I_{1}}(u)+\sum\limits_{v\in I_{2}}d^{\geq}_{I_{1}\setminus I_{2}}(v)\leq 5|I_{1}\triangle I_{2}|.$
3.4 Proof of Theorem 1.4
We prove that disk graphs have an $11$-simplicial vertex, i.e. the set of its neighbors is the union of $11$ cliques.
For a disk graph $G$ and $v\in V(G)$, denote by $D_{G}(v)$ the disk corresponding to $v$ and $\rho_{G}(v)$ the radius of $D_{G}(v)$. Let $v_{0}\in V(G)$ be a vertex with minimum radius $\rho_{G}(v_{0})$ among all vertices in $G$. By rescaling, we may assume $\rho(v_{0})=1$. We shall prove that $v_{0}$ is the desired $11$-simplicial vertex.
For each neighbor $v\in N_{G}(v)$, define a new disk $D^{\prime}(v)$ with radius 1 internally tangent to $D_{G}(v)$ such that the common point $A=D^{\prime}(v)\cap D_{G}(v)$ lies on the line going through centers of $D_{G}(v_{0})$ and $D_{G}(v)$ (see Fig. 3.2). As each new disk $D^{\prime}(v)$ is contained in the original $D_{G}(v)$, for any $v,w\in N_{G}(v_{0})$, if $D^{\prime}(v)\cap D^{\prime}(w)\neq\varnothing$, then $vw\in E(G)$. As depicted in Fig. 3.2, the family of all new disks ${\mathcal{D}}=\{D^{\prime}(v):v\in N_{G}(v_{0})\}$, all of which intersect $D_{G}(v_{0})$, can be pierced by at most 11 points, implying that $v_{0}$ is an 11-simplicial vertex in $G$.
[…]
Figure 3.1: Create the disk $D^{\prime}$ according to the disks $D_{0}$ and $D$
Figure 3.2: Any unit disk $D$ which intersects $D_{0}$ must contain one of the eleven points
4 Circle graphs: Proof of Theorem 1.5
Consider a circle $O$, where the vertex set of graph $G$ consists of $n$ distinct chords in $O$. By perturbation, we may assume that none of these $n$ chords are diameters of $O$. For each $v\in V(G)$, let $C_{v}$ denote its corresponding chord and $A_{v}$ the minor arc associated to $C_{v}$.
For two non-crossing chords $C_{a},C_{b}$, we say that $C_{a}$ covers $C_{b}$ if $A_{a}\supseteq A_{b}$. Furthermore, we say that $C_{a}$ directly covers $C_{b}$ if there does not exist any other chord $C_{c}$ with $A_{a}\supseteq A_{c}\supseteq A_{b}$ (see Fig. 4.2). For $X\in\mathcal{I}_{\textup{max}}(G)$ and $x\in X$, let $K_{X}(x)$ be the set of vertices in $X$ which are covered by $x$ and let $k_{X}(x):=|K_{X}(x)|$. Obviously, $10\leq k_{X}(x)\leq cn-$ 1. Note also that for any $a,b\in X$, if $C_{a}$ covers $C_{b}$, then $k_{X}(a)>k_{X}(b)$. Now, for each $I\in\mathcal{I}_{\textup{max}}(G)$ and $v\in I$, assign it the color $k_{I}(v)$.
ABCD
Figure 4.1: $A$ covers $B,C,D$, but $A$ does not directly cover $C,D$
Figure 4.2: Any chord can cross at most two distinct chords in $Z_{I}$
Claim 4.1.
Each vertex in $V(G)$ receives at most one color.
Proof of claim.
Suppose that some vertex $u$ receives two distinct colors, say for $X,Y\in\mathcal{I}_{\textup{max}}(G)$, $k_{X}(u)=i<j=k_{Y}(u)$. Note that chords in $X\setminus K_{X}(u)$ are disjoint from those in $K_{Y}(u)$ as they lie in different sides of $C_{u}$. Therefore, $(X\setminus K_{X}(u))\cup K_{Y}(u)$ is also an independent set of size $cn-i+j>cn$, a contradiction. ∎
Next, for $I\in\mathcal{I}_{\textup{max}}(G)$, consider a vertex $v\in I$ with minimum $k_{I}(v)$. The minimality of $v$ implies that it does not cover any other vertex in $I$ and so $k_{I}(v)=0$. Thus, we have the following claim.
Claim 4.2.
For each $I\in\mathcal{I}_{\textup{max}}(G)$, the set $Z_{I}=\{v\in I:k_{I}(v)=0\}\neq\emptyset$.
Note that for any $I\in\mathcal{I}_{\textup{max}}(G)$ and any three pairwise non-crossing chords in $Z_{I}$, any other chord can cross at most two of them (see Fig. 4.2). Thus, the number of edges between $Z_{I}$ and $V(G)\setminus I$ is at most $2(1-c)n$ and there exists a vertex $v\in Z_{I}$ with $d_{G}(v)\leq\frac{2(1-c)n}{|Z_{I}|}$. Then by 2.1, $h(G)\leq\frac{2(1-c)n}{|Z_{I}|}+1$. Consequently, if there is $I$ with $|Z_{I}|>\frac{\log{n}}{\log\log{n}}$, then we are done. We may then assume that $|Z_{I}|\leq M:=\frac{\log{n}}{\log\log{n}}$ for all $I\in\mathcal{I}_{\textup{max}}(G)$.
Let $R\subseteq V(G)$ denote the set of vertices that do not belong to any maximum independent set. By 4.1, every vertex in $V(G)\setminus R$ is assigned a unique color in $\{0,1,\ldots,cn-1\}$. Partition $V(G)\setminus R$ by colors $V(G)\setminus R=V_{0}\cup V_{1}\cup V_{2}\cup\cdots\cup V_{cn-1}$, where any vertex $v\in V_{i}$ satisfies $k_{I}(v)=i$ for any $I\in\mathcal{I}_{\textup{max}}(G)$. By 4.2, $V_{0}$ is a hitting set for $\mathcal{I}_{\textup{max}}(G)$. However, it is hard to bound the size $V_{0}$ directly. We will instead use multiple layers of $V_{i}$ instead.
Fix an $I\in\mathcal{I}_{\textup{max}}(G)$ with $|Z_{I}|\leq M$.
Claim 4.3.
For any $v\in I$, $v$ directly covers at most $M$ other vertices in $I$.
Proof of claim.
Let $U\subseteq I$ be the set of vertices in $I$ that are directly covered by $v$. Then no vertex in $U$ covers another. This, together with the fact that $U\subseteq I$ is independent, infers that the minor arcs $A_{u}$, $u\in U$, are pairwise disjoint. Then, considering a vertex $u^{\prime}$ with minimal color with $A_{u^{\prime}}\subseteq A_{u}$ for each $u\in U$, we see that $|Z_{I}|\geq|U|$ finishing the proof. ∎
Claim 4.4.
$I\cap\bigcup_{i=1}^{M}V_{i}\neq\varnothing$ and $|I\cap V_{i}|\leq M$ for all $i\in[cn-1]$. In particular, $h(G)\leq|\bigcup_{i=1}^{M}V_{i}|$.
Proof of claim.
Consider a vertex $u\in I\setminus Z_{I}$ with minimum color. By the minimality, $u$ can cover only vertices in $Z_{I}$ and so $k_{I}(u)\leq|Z_{I}|$.
Fix $i\in[cn-1]$, then the minor arcs $A_{u}$, $u\in I\cap V_{i}$, are pairwise disjoint. Again, considering a vertex $u^{\prime}$ with minimal color with $A_{u^{\prime}}\subseteq A_{u}$ for each $u\in I\cap V_{i}$, we see that $|Z_{I}|\geq|I\cap V_{i}|$. ∎
Let $t$ be the largest positive integer such that $(2M)^{t+1}\leq\lfloor\frac{cn}{M+1}\rfloor$. Then by 4.4, we see that $I$ contains vertices with color at least $(2M)^{t+1}+1$, that is, $I\setminus(\bigcup_{i=1}^{(2M)^{t+1}}V_{i})\neq\varnothing$.
Claim 4.5.
For each $0\leq j\leq t$, $I\cap\bigcup_{k=(2M)^{j}+1}^{(2M)^{j+1}}V_{k}\neq\varnothing$. In particular, $h(G)\leq|\bigcup_{k=(2M)^{j}+1}^{(2M)^{j+1}}V_{k}|$.
Proof of claim.
We prove it by induction on $0\leq j\leq t$. The base case $j=0$ follows from 4.4. For the inductive step, suppose the claim holds for all $0\leq j^{\prime}<j\leq t$. Consider a vertex $w$ in $I\setminus(\bigcup_{i=1}^{(2M)^{j+1}}V_{i})$ with minimum color. Then, as $w$ can directly cover at most $M$ vertices in $I$ by 4.3, there must exist a vertex $u$ covering at least $\frac{(2M)^{j+1}+1-M}{M}\geq(2M)^{j}+1$ vertices. On the other hand $u$ covers at most $(2M)^{j+1}$ vertices by the minimality of $w$. This finishes the proof. ∎
As each $\bigcup_{k=(2M)^{j}+1}^{(2M)^{j+1}}V_{k}$, $j=0,1,\ldots,t$, is a hitting set for $\mathcal{I}_{\textup{max}}(G)$, by averaging, there is a hitting set of size at most $\lfloor\frac{n}{t+1}\rfloor$. By the choice of $t$, $(2M)^{t+2}\geq\frac{cn}{M+1}$ and so $t\geq\frac{\log\frac{cn}{M+1}}{\log(2M)}-2\geq\frac{\log n}{4\log\log n}$. Thus, $h(G)\leq\frac{4n\log\log{n}}{\log{n}}$. The proof is finished.
5 Concluding remarks
In conclusion, our study delves into the longstanding open problem proposed by Bollobás, Erdős and Tuza, which posits that for any $n$-vertex graph with $\alpha(G)=\Omega(n)$, it suffices to employ $o(n)$ vertices to hit all maximum independent sets. Our investigation draws inspiration from the remarkable result, along with its elegant proof, by Alon [3], demonstrating validity of 1 for regular graphs with large independence number, as well as the work of Hajebi, Li, and Spirkl [35], establishing the existence of small hitting sets for induced $P_{5}$-free graphs.
Our primary contribution lies not only in confirming the validity of 1 across several significant graph classes but also in synthesizing concepts from diverse fields to either construct or demonstrate the existence of small hitting sets. We envision our ideas serving as a springboard for further exploration aimed at tackling this conjecture.
It is intriguing to note that the conjecture remains open even when $\alpha(G)=(\frac{1}{2}-\varepsilon)|G|$ for any small $\varepsilon>0$. We believe that addressing this weaker conjecture presents its own set of challenges worthy of pursuit.
5.1 Constant-sized hitting sets in various sporadic graphs
In addition to the graph classes we have discussed in Section 2, in many other classes of geometric graphs, we can also prove the existence of hitting sets of constant size. We select some representative ones to introduce the sketches briefly. For a configuration $S$ in high dimensional space $\mathbb{R}^{d}$, let $\mu(S)$ be the Lebesgue measure of $S$, and $|\bm{x}-\bm{y}|$ represent the Euclidean distance between points $\bm{x},\bm{y}\in\mathbb{R}^{d}$, we define the diameter of $S$ as $\textup{Diam}(S):=\sup\{|\bm{a}-\bm{b}|:\bm{a},\bm{b}\in S\}$. Observe that in an intersection graph of configurations with positive measures and bounded diameter $D$, there is no large induced star, because the ball of radius $2D$ can only contain bounded number of pairwise disjoint configurations. Therefore, the following result follows.
Proposition 5.1.
Let $\varepsilon>0$ be a real number and $d\geq 1$ be a positive integer. Let $\mathcal{S}:=\{S_{1},S_{2},\ldots,S_{n}\}\subseteq\mathbb{R}^{d}$ be a collection of $n$ sets such that for each $i\in[n]$, $\mu(S_{i})>\varepsilon$ and $\textup{Diam}(S_{i})\leq D$ for some constant $D>0$. Let $G$ be the intersection graph of the sets in $\mathcal{S}$, then $h(G)=O_{\varepsilon,D,d}(\frac{n}{\alpha(G)})$. In particular, if $\alpha(G)=cn$ for some $c>0$, then $h(G)=O_{\varepsilon,D,d,c}(1)$.
In low dimensional space, the interval graph is formed from a set of intervals on the real line, with a vertex for each interval and an edge between vertices whose intervals intersect. It was known that any interval graph is an incomparability graph, therefore, by 1.7, we can see $h(G)\leq\frac{n}{\alpha(G)}$ when $G$ is an $n$-vertex interval graph.
An interesting generalization of interval graph is circular-arc graph, which is the intersection graph of a set of arcs on the circle. It has one vertex for each arc in the set, and an edge between every pair of vertices corresponding to arcs that intersect. Observe that, for any given maximum independent set $I$, if a vertex $v\in V(G)\setminus I$ such that $|N_{G}(v)\cap I|\geq 4$, then $v$ cannot belong to any maximum independent set in $G$. Based on this, the following result follows.
Proposition 5.2.
Let $G$ be a circular-arc graph, then we have $h(G)\leq\frac{3n}{\alpha(G)}-2$, and in particular if $\alpha(G)=cn$ for some $c>0$, we have $h(G)\leq\frac{3}{c}-2$.
5.2 Extension of circle graphs
We can also extend the technique in the proof of Theorem 1.5 to a more general class of graphs. More precisely, a circle $k$-gon is the region that lies between $k$ or fewer non-crossing chords of a circle, no chord connecting the arcs between two other chords. The sides of a circle $k$-gon are either chords or arcs of the circle. A graph is a circle k-gon graph if it is the intersection graph of circle $k$-gons in a common circle. One can see that any circle graph is indeed a circle $1$-gon graph. Using the almost identical argument, we can show the following more general result.
Theorem 5.3.
Let $k\geq 1$ be a positive integer and let $G$ be a circle $k$-gon graph with $\alpha(G)=\Omega(n)$, then we have $h(G)\leq O(\frac{n\log\log{n}}{\log{n}})$.
5.3 The bridge between the balanced separators and the hitting sets
For an $n$-vertex graph $G$, a subset $S\subseteq V(G)$ is a balanced separator if every component of $G\setminus X$ has at most $\frac{2n}{3}$ vertices. Equivalently, $S\subseteq V(G)$ is a balanced separator if there is a partition $V:=S\cup V_{1}\cup V_{2}$ with $|V_{1}|,|V_{2}|\leq\frac{2|V(G)|}{3}$ such that no vertex in $V_{1}$ is adjacent to any vertex in $V_{2}$.
Here, we establish a straightforward connection between the balanced separator and the hitting set. Let $\mathcal{G}_{\varepsilon}$ be a hereditary family of graphs such that for every graph $G\in\mathcal{G}$, there exists a balanced separator of size at most $\varepsilon|V(G)|$. Note that for any $n$-vertex graph, we can repeatedly take balanced separators and consider the subgraphs induced by smaller components by 2.1. Note that the union of those separators forms a hitting set, which is of size at most $\sum\limits_{i}\frac{\varepsilon n}{2^{i-1}}\leq 2\varepsilon n$.
Proposition 5.4.
Let $G\in\mathcal{G}_{\varepsilon}$ be an $n$-vertex graph, then $h(G)\leq 2\varepsilon n$.
Therefore, for those hereditary graphs having sublinear balanced separators, they also have sublinear hitting sets. For instance, a classical result of Lipton and Tarjan [45] shows that every $n$-vertex planar graph contains a balanced separator of size $O(\sqrt{n})$. This result on planar graphs has also been generalized to graphs embedded in a surface of bounded genus [33], graphs with a forbidden minor [4], and intersection graphs of some geometric objects [23, 46].
In addition to the aforementioned classical results, a recent study by Dvořák, McCarty, and Norin [18] provided a sufficient condition for an intersection graph of compact convex sets in $\mathbb{R}^{d}$ to possess a balanced separator of sublinear size. Furthermore, Dvořák and Wood [19] have elucidated the connection between balanced separators and other parameters such as tree-width, path-width, and tree-depth. For additional hereditary families of graphs with sublinear balanced separators, we recommend interested readers to refer to [18] and the references therein.
5.4 Some open problems
5.4.1 Graphs with bounded VC-dimension
Theorem 1.8 indicates that any graph with VC-dimension one has a very small hitting set. While for any larger positive integer $d\geq 2$, it remains an interesting open problem.
Conjecture 2.
Let $d\geq 2$ be a positive integer, and let $G$ be an $n$-vertex graph with VC-dimension at most $d$. If $\alpha(G)=\Omega(n)$, then $h(G)=o(n)$.
We anticipate that insights from several papers [1, 25, 26, 27, 47, 51] may prove instrumental in our exploration. Motivated by the recent work of Nguyen, Scott and Seymour [47], who proved that graphs with bounded VC-dimension have Erdős-Hajnal property, one might expect that $h(G)\leq\textup{Poly}(\omega(G))$ holds for the graphs with bounded VC-dimension. However, this is not true. Alon’s construction in [3, Theorem 1.3] showed that there is a graph $G$ with bounded VC-dimension and small clique number, but $h(G)$ cannot be upper bounded by any function of $\omega(G)$.
Proof of Theorem 1.1.
Let $k$ be a positive integer and $K:=\{1,2,\ldots,2k\}$. The set of vertices of $G$ is the set of all ordered pairs $(i,j)$ with distinct $i,j\in K$. Note that the number of vertices is $n=2k(2k-1)$. A pair of vertices $(a,b)$ and $(c,d)$ are adjacent if $b=c$ or $a=d$. Alon [3] has already proved that $h(G)\geq\frac{\sqrt{n}}{2}$. It then remains to show $\omega(G)=3$ and VC-dimension is at most $3$. First, it is easy to check that any triangle in $G$ must be of the form $(a,b),(b,c),(c,a)$, where $a,b,c\in K$ are distinct, moreover, it is obvious that these three vertices have no common neighbor. For the VC-dimension of $G$, suppose $S=\{(a_{i},b_{i}):i\in[4]\}$ is a shattered set. Let $(x,y)$ be the vertex such that $N_{G}((x,y))\supseteq N_{G}(S)$. Then for each $1\leq i\leq 4$, either $a_{i}=y$ or $b_{i}=x$. If for each $1\leq i\leq 4$, $a_{i}=y$ (or $b_{i}=x$ respectively), then there is no vertex $(z,w)$ such that $|N_{G}((z,w))\cap S|=3$. Therefore, without loss of generality, we can assume there exist three vertices in $S$ of the form $(a_{1},x),(a_{2},x),(y,b_{3})$ (or $(b_{1},y),(b_{2},y),(a_{3},x)$), then it is easy to check that these three vertices have no common neighbor other than $(x,y)$. This finishes the proof. ∎
5.4.2 String graphs
In a fundamental sense, curves represent the most versatile geometric objects on the plane, defined as the images of continuous functions $\phi:[0,1]\rightarrow\mathbb{R}^{2}$. A string graph emerges as the intersection graph of such curves. Notice that the majority of graph families examined within this paper align with the concept of string graphs. Consequently, the following conjecture assumes significant relevance.
Conjecture 3.
Let $G$ be an $n$-vertex string graph.
  1. (1)
    If $\alpha(G)=\Omega(n)$, then $h(G)=o(n)$.
  2. (2)
    There exists some constant $s>0$ such that $h(G)\leq\omega(G)^{s}$.
The significant finding by Fox and Pach [24] might be helpful, which asserts that any dense string graph must contain a dense incomparability graph as a spanning subgraph. 3(2) is also motivated by a recent breakthrough by Tomon [53], who demonstrated that string graphs have the Erdős-Hajnal property (also see a more flexible result [28]). Building on the relationship discovered in [35], a proof of 3(2) would offer a novel proof of Tomon’s theorem and possibly lead to potential quantitative improvements.
We are also interested in the family of interval filament graphs introduced by Gavril [30]. A graph is an interval filament graph if it is the intersection graph of continuous non-negative functions defined on closed intervals with zero-values on their endpoints. One of the motivations is that the family of interval filament graphs is a common generalization of the incomparability graphs, the circle graphs and the chordal graphs.
Conjecture 4.
Let $G$ be an $n$-vertex interval filament graph with $\alpha(G)=\Omega(n)$, then $h(G)=o(n)$.
5.4.3 Optimal construction
We believe that 1 is true, but currently we are unable to prove it. Conversely, we conjecture that the $o(n)$ term in the conjecture is optimal, suggesting that the following conjecture may be valid.
Conjecture 5.
For any positive constant $\varepsilon>0$, there exists an $n$-vertex graph $G$ with $\alpha(G)=cn$ for some constant $c>0$ and $h(G)\geq n^{1-\varepsilon}$.
Acknowledgement
Zixiang Xu would like to express sincere gratitude to Prof. Hong Liu for his extensive suggestions on the writing of this draft, and for providing the proof of 1.6. The authors also thank Prof. Noga Alon for helpful comments.